A roller derby car with a mass of 120kg starts at the top of the hill (20 meters high) how fast is the car going at point B with a height of 15 meters? (use g = 10m/s/s and create a LOL to help solve the problem)
1
Expert's answer
2020-05-28T13:14:15-0400

Apply energy conservation principle: the potential energy at height H=20 m converts into kinetic energy and potential energy h=15 m:


"mgH=mgh+\frac{mv^2}{2},\\
v=\sqrt{2g(H-h)}=\sqrt{2\cdot9.8(20-15)}=9.9\text{ m/s}."